# Deep Generative Models for Detector Signature Simulation: An Analytical Taxonomy

### v1

@@ 4 Applications
In this section, we give an overview of the applications of DGMs in detector signature simulation. Specifically, we categorize them into three concrete branches: statistics amplification, Amortised generation, and OOD simulation. Then, we illustrate their formulations in detector response generation and how different approaches and incorporation of inductive biases lead to success in various applications. “Inductive biases” in DGMs refer to the assumptions the model makes to predict outputs for both in- and out-of-distribution inputs. These biases are inherent in the model’s architecture and learning algorithm such as symmetries and adversarial robustness that guide the model’s learning process [214, 215, 216, 217, 218, 219, 220, 221, 45, 222]. In general, for DGMs in particle physics, the inductive biases could include distributional assumptions such as smoothness inductive bias, structural assumptions such as relational inductive bias or geometry awareness, and Physics-informed Assumptions such as Energy-Momentum conservation bias.
@@ 4.1 Statistics Amplification
As surrogate models are being used for detector signature emulation to do sample amplification [223], it is very important to quantify the statistical power of the generated dataset. Sample amplification is a procedure where there exists a map that takes a finite, initial subset of data and generates an extended one. In this regime, DGMs work as classical parametric fits [224, 225] to the training that also works as data augmentation methods. The question then would be to which extent they can interpolate. In other words, how can one quantify the diversity and uncertainty of DGMs? This involves understanding the limitations and potential biases of the DGMs and possibly introducing methods to quantify the diversity or amplification factors [226, 227]. For amplification of statistics, the smoothness assumption where the physics probability densities are smooth is very natural. In DGMs, the smoothness inductive bias is an assumption that similar inputs will produce similar outputs up to model interpolation. Regularization techniques in machine learning, which are used to encourage the model to learn smoother functions, are a common manifestation of the smoothness inductive bias.
The smoothness inductive bias is studied in [228, 11, 12] for a simplified VAE-GAN [84]. Their study revealed that individual-generated shower samples contain less information compared to a single real data point. When the number of generated shower samples increases significantly, the information contained within the generated sample set eventually levels off. As a result, they demonstrate the ability of DGMs to not only sample from implicitly defined distributions but also to leverage enhanced interpolation or fitting capabilities. Similar results are being achieved in [14] from an Information Theoretic perspective while introducing a versatile diversity measure for detector signature simulation.
To improve interpolation, it’s crucial to consider how to diversify samples. In their work, Hashemi et al. [45] enhance the diversity and complexity [222, 229] of generated samples by introducing Self-Supervised Learning [230] techniques that can be incorporated in any DGMs. In particular, they introduce a Uniformity loss that helps the model to avoid mode collapse and to generate more diverse samples by maintaining the “uniformity of information” inductive bias for the discriminator. Kansal et al. [231] study in detail the quantifying of the diversity of the generated samples by analyzing various metrics and measures for the jet point cloud generation task. They show that the message-passing GAN [150] performs better than set-transformer [232].
@@ 4.2 Amortised Generation
In an amortized generation, the sub-tasks are either “Fast Simulation” to overcome the high time complexity of detector simulation or “Data Compression” [78, 121, 110, 233, 119] to overcome the high space complexity of detector simulation. Based on the different types of data manifolds, various inductive biases are being incorporated. One can categorize different representations into two main classes, “geometry-dependent” and “geometry-independent” representations. Geometry-dependent approaches look at the detector signatures as grid-like structures. For example, 1d fixed representations come with sequential local and translation invariance [234, 235]. 2D representation with local and translation-invariant inductive bias, which was studied for showers, jets, and tracks in depth for different detectors. Then, in order to capture the detector’s layer-by-layer association and correspondence, 3d grid-based models [132, 133, 62, 61, 82] are studied. However, as a consequence of translation invariance, it has the drawback of stationary assumption over the temporal/spatial features.
@@
Geometry-independent approaches, on the other hand, are more suitable for simulating detector signatures due to the inherent variable length of data and their heterogeneity and sparsity. For example, graph-based models [103, 149, 107, 150, 59], besides the variable-length assumption and relational inductive bias, it assumes graph isomorphism inductive bias as well. This bias ensures that the model focuses on the structural information contained in the graph rather than the specific labeling of nodes. Another geometry-independent approach is considering the set representation of detector signatures. The most important property with set-based models are the permutation equivariant encoders for jet, shower or track constituents [211, 173, 190, 188, 86, 14] or sensors [45], and possibly permutation invariant loss functions [107, 59, 231]. For normalizing flows, equivariance under permutation group action is more non-trivial. Proved by Köhler et al. [236], given a Flow-based model $F$ such that the set creation yields an exchangeable distribution, the update is permutation equivariant and invertible, and $p_{\theta}$ denotes the model likelihood, then $(F,-\log p_{\theta})$ is permutation equivariant.
@@
While maintaining the variable-length assumption, dropping the permutation equivariance would correspond to a sequential inductive bias of detector responses either layer-by-layer [14, 213], or hit-by-hit [207, 208, 209]. While autoregressive models show a better predictive capability [194], the sequential nature of these models can be a disadvantage when it comes to high-dimensional computation, as the sampling is rather slow.
@@ 4.3 Out-Of-Distribution (OOD) Generation
DGMs for OOD and zero-shot learning is an exciting area of research that holds significant potential across various fields [237, 238, 239, 240, 241, 242, 243, 244] including, but not limited to, Drug Discovery, material design, and weather forecasting. Traditional methods for the synthetic generation of objects with enhanced or specific properties are often iterative and costly, requiring extensive manual work or heavy computational resources.
@@
In contrast, DGMs with zero-shot capability can deal with new scenarios that are not explicitly present in the training data, making them highly desirable in a wide range of problems. OOD generation of detector signatures is an emerging field in High Energy Physics (HEP) as well [68, 158, 245, 176, 209, 14]. For example, for real PXD data at Belle II [36], an important challenge in working is its reliance on actual experiments for gathering the necessary random triggers. Consequently, real luminosity and beam-parameter-dependent PXD background data that go beyond current experimental limits are unavailable, leaving us reliant on computationally demanding simulations. This underscores the critical need for a surrogate model that can effectively generalize to OOD luminosity regions. The current main challenge remains to be the optimization of DGMs, based on available information while avoiding overfitting, and the generalization to cases in which information is scarce or altogether absent, such as extrapolation to beam parameters, energies, luminosities, and geometries where there are no data.
@@
The first example of such OOD detector response generation is the CaloGAN paper [68], where they very briefly show they can generate showers beyond the training incident energy conditions. DijetGAN [158], also discusses extrapolation to new Beyond the Standard Model (BSM)-dependent OOD regions of dijet invariant mass. Anderlini et al. [176] evaluates the uncertainty of GANs in new momentum regions through Background efficiency comparison. Their analysis shows that for Kaons, where the background efficiency does not decrease linearly in the OOD, high-momentum phase-space regions, their ensemble model fails to capture this behavior. Whereas, for Muons, due to the monotonic behavior of the background efficiency, the extrapolation shows better results.
@@
A very important inductive bias for OOD generation is the variable-length assumption, especially in the context of detector signature simulation, due to the inherently variable nature of the data, characterized by a fluctuating number of detector responses in individual events. DGMs with a variable-length inductive bias have a promising capability to generate sets/sequences of differing lengths, contingent on the inherent complexity or dynamism of the event being generated. This equips the model with better adaptability when faced with novel and unseen scenarios. Liu et al. [208, 209], by using an ARM approach (length-independent), tries to extrapolate to unseen detector geometries where the extrapolation domain is calorimeter cell sizes. However, they show that the high granularity is a bottleneck to their approach. Hashemi et al. [14] address the challenge of OOD simulation for high granular and sparse events (variable intra-event cardinality) by introducing YonedaVAE that can achieve both length and context extrapolation. Their model is trained on random trigger PXD background data from an early experiment of Belle II with a peak recorded luminosity of $1.42\times 10^{34}\text{cm}^{-2}\text{s}^{-1}$ and a mean occupancy of 0.06% ($\mathcal{O}(100)$ hit multiplicity). It was then tested on the previously unseen data from a later experiment of Belle II, which had nearly double the peak luminosity — $2.68\times 10^{34}\text{cm}^{-2}\text{s}^{-1}$ — and a mean occupancy of 0.32% ($\mathcal{O}(10^{5})$ hit multiplicity). The length extrapolation here was when the model had to infer how an event with a higher (than training) cardinality looks like. Context extrapolation means that the model has to produce the correct distribution of sparse cardinalities over all sensors, given only the maximum cardinality per event.
@@ 5 Outlook: Challenges and Opportunities
Physics-Informed Generative Models. One of the most pressing challenges in the field of deep generative models for simulating detector signatures in particle physics is the incorporation of physics-based constraints or rules into the model architecture. While current methodologies often focus on data-driven approaches (permutation equivariance is a special case), they sometimes overlook the underlying physics and symmetries that govern the data. This could lead to models that are statistically accurate but physically implausible. Future work could focus on models that not only learn from data but also respect the governing symmetries [221], thereby ensuring both statistical and physical fidelity in the generated signatures.
@@
Precision and Uncertainty Quantification. One of the less explored but critically important areas in the application of deep generative models for simulating high-granular detector signatures is the quantification of model uncertainty and the evaluation of precision [13]. While these models can generate data that statistically resemble the target data, the question of how accurately they capture the underlying physical processes remains. Traditional metrics [178, 231, 45] like log-likelihood or NN-based metrics provide some measure of the model’s performance but often fall short in capturing the nuances of physical validity and uncertainty. Moreover, the stochastic nature of these models introduces an inherent level of uncertainty in the predictions, which needs to be rigorously quantified to make the models truly useful. Future work could focus on the development of new evaluation metrics and techniques that not only assess the model’s ability to replicate known phenomena but also its reliability in predicting new high-granular detector signatures. This could involve Bayesian approaches [246, 247, 248], ensemble methods [176], or even hybrid models [249] to provide error bars along with predictions.
@@
Computational Efficiency and Real-Time Optimization. The increasing complexity of particle physics experiments, characterized by high event rates and pile-up conditions, poses significant challenges for real-time data analysis and decision-making for ultra-high granular detectors. Surrogate models, as discussed in this paper, provide very affordable solutions; however, optimizing their real-time performance is still a bottleneck. One promising avenue is the use of Federated Learning (FL) [250], which allows for decentralized optimization across multiple nodes, thereby leveraging the computational resources of the entire experimental setup. FL enables models to train on large real-time data sets and reduce biases associated with locally trained models. Meta-learning techniques [251] are also strong candidates to enable models to quickly adapt to new experimental data or detector conditions, reducing the time required for retraining. Recently, the MetaHEP [252] project embarked on this path. Additionally, another promising avenue for future research lies in the direction of differentiable programming (DP) [6, 253, 254, 255]. Through DP, software becomes differentiable through automatic differentiation (AD) [256], enabling efficient gradient computation to understand the influence of input variations on output predictions. Such gradient-based insights could be invaluable for various downstream tasks. Utilizing DP frameworks would allow particle physics simulation tools to be integrated into machine learning pipelines in an end-to-end manner, thereby facilitating joint optimization for enhanced computational efficiency.
@@
Quantum Generative Models. The advent of quantum computing offers a new avenue for tackling the computational challenges in particle physics simulations [257]. Quantum Generative Models (QGM) could potentially revolutionize the field by providing exponential speed-ups for certain types of problems. However, the practical implementation of QGMs for particle physics is still in its infancy [170, 258, 259, 260], and significant challenges related to error correction, qubit stability, and algorithmic design remain. Future research could focus on the development and validation of QGM architectures specifically tailored for detector signature simulations.
@@
Extrapolation Beyond Training Data. One of the limitations of current generative models is their ability to generalize beyond the scope of the training data. Given the high-dimensional nature of HEP data, the inevitability of sparse datasets arises due to limited sample sizes, leading us into zones with inadequate training data. If the model demonstrates strong extrapolation capabilities, it doesn’t just fill gaps in areas completely lacking data; it also provides a more reliable interpretation of regions where data is minimal. This is particularly crucial in particle physics, where there is a need to explore uncharted territories of parameter space [245], kinematic regions [7, 176], luminosities [14] or detector geometries [209]. Developing models capable of reliable extrapolation is an open challenge and involves techniques like uncertainty quantification and the incorporation of prior physical knowledge and inductive bias into the model.
@@
Scalability and Ultra-High Granularity Challenges. The next generation of particle detectors will feature ultra-high granularities, as we defined in Section 3.1.3, leading to an explosion in the dimensionality of the data [261], for instance, the high Granularity Calorimeter (HGCAL) [33] at CMS [64] with roughly $6$ million channels, the ultra-high resolution PXD at Belle II [36] with more than $7.5$ million-pixel channels, or the EPICAL-2 [65] electromagnetic calorimeter with $12.5$ million-pixel channels. Except for some efforts [74, 14], current generative models struggle to handle such ultra-high-dimensional data efficiently. Even these efforts have not delved deeply into the issue of uncertainty quantification. This gap underscores the urgent need for future work that not only addresses the development of scalable algorithms and architectures that can handle ultra-high granularity detectors with irregular geometries but also rigorously quantifies the uncertainties to ensure the physical validity and reliability of the generated detector signatures.

### v2

@@ 4 Applications
In this section, we give an overview of the applications of DGMs in detector signature simulation. Specifically, we categorize them into four concrete branches: statistics amplification, Amortised generation, OOD simulation, and anomaly detection. Then, we illustrate their formulations in detector response generation and how different approaches and incorporation of inductive biases lead to success in various applications. “Inductive biases” in DGMs refer to the assumptions the model makes to predict outputs for both in- and out-of-distribution inputs. These biases are inherent in the model’s architecture and learning algorithm such as symmetries and adversarial robustness that guide the model’s learning process [zhao_bias_2018, battaglia_relational_2018, zietlow_demystifying_2021, krippendorf_detecting_2020, dillon_symmetries_2021, barenboim_symmetry_2021, tombs_method_2022, desai_symmetrygan_2022, hashemi_ultra-high-resolution_2023, haochen_theoretical_2023]. In general, for DGMs in particle physics, the inductive biases could include distributional assumptions such as smoothness inductive bias, structural assumptions such as relational inductive bias or geometry awareness, and Physics-informed Assumptions such as Energy–Momentum conservation bias.
@@ 4.1 Statistics Amplification
As surrogate models are being used for detector signature emulation to do sample amplification [axelrod_sample_2019], it is very important to quantify the statistical power of the generated dataset. Sample amplification is a procedure where there exists a map that takes a finite, initial subset of data and generates an extended one. In this regime, DGMs work as classical parametric fits [carrazza_compressing_2021, chahrour_comparing_2022] to the training that also works as data augmentation methods. The question then would be to which extent they can interpolate. In other words, how can one quantify the diversity and uncertainty of DGMs? This involves understanding the limitations and potential biases of the DGMs and possibly introducing methods to quantify the diversity or amplification factors [hao_data_2019, axelrod_statistical_2022]. For amplification of statistics, the smoothness assumption where the physics probability densities are smooth is very natural. In DGMs, the smoothness inductive bias is an assumption that similar inputs will produce similar outputs up to model interpolation. Regularization techniques in machine learning, which are used to encourage the model to learn smoother functions, are a common manifestation of the smoothness inductive bias.
The smoothness inductive bias is studied in [butter_amplifying_nodate, butter_ganplifying_2021, bieringer_calomplification_2022] for a simplified VAE-GAN [buhmann_getting_2021]. Their study revealed that generated shower samples contain less information compared to a single real data point in the low sample size regime. When the number of generated shower samples increases significantly, the information contained within the generated sample set eventually levels off. As a result, they demonstrate the ability of DGMs to not only sample from implicitly defined distributions but also to leverage enhanced interpolation or fitting capabilities. Similar results are being achieved in [hashemi_deep_2024] from an Information Theoretic perspective while introducing a versatile diversity measure for detector signature simulation.
To improve interpolation, it is crucial to consider how to diversify samples. In their work, Hashemi et al. [hashemi_ultra-high-resolution_2023] enhance the diversity and complexity [haochen_theoretical_2023, shwartz-ziv_compress_2023] of generated samples by introducing Self-Supervised Learning [balestriero_cookbook_2023] techniques that can be incorporated in any DGMs. In particular, they introduce a Uniformity loss that helps the model to avoid mode collapse and to generate more diverse samples by maintaining the “uniformity of information” inductive bias for the discriminator. Kansal et al. [kansal_evaluating_2023] study in detail the quantifying of the diversity of the generated samples by analyzing various metrics and measures for the jet point cloud generation task. They show that the message-passing GAN [kansal_particle_2022] performs better than set-transformer [lee_set_2019].
@@ 4.2 Amortised Generation
In an amortized generation, the sub-tasks are either “Fast Simulation” to overcome the high time complexity of detector simulation or “Data Compression” [srebre_generation_2020, huang_efficient_2021, collins_machine-learning_2022, bengtsson_baler_2023, huang_fast_2023] to overcome the high space complexity of detector simulation. Based on the different types of data manifolds, various inductive biases are being incorporated. One can categorize different representations into two main classes, “geometry-dependent” and “geometry-independent” representations. Geometry-dependent approaches look at the detector signatures as grid-like structures. For example, one-dimensional fixed representations come with sequential local and translation invariance [mu_photon_2021, regadio_synthesis_2022]. Two-dimensional, 2D, representation come with local and translation-invariant inductive bias, which was studied for showers, jets, and tracks in depth for different detectors. Then, in order to capture the detector’s layer-by-layer association and correspondence, 3d grid-based models [khattak_three_2018, vallecorsa_3d_2019, belayneh_calorimetry_2020, khattak_fast_2021, mikuni_score-based_2022] are studied. However, as a consequence of translation invariance, it has the drawback of stationary assumption over the temporal/spatial features.
@@
Geometry-independent approaches, on the other hand, are more suitable for simulating detector signatures due to the inherent variable length of data and their heterogeneity and sparsity. For example, graph-based models [hariri_graph_2021, kansal_graph_2021, tsan_particle_2021, kansal_particle_2022, di_bello_conditional_2022], besides the variable-length assumption and relational inductive bias, it assumes graph isomorphism inductive bias as well. This bias ensures that the model focuses on the structural information contained in the graph rather than the specific labeling of nodes. Another geometry-independent approach is considering the set representation of detector signatures. The most important property with set-based models are the permutation equivariant encoders for jet, shower or track constituents [finke_learning_2023, buhmann_epic-gan_2023, leigh_pc-jedi_2023, mikuni_fast_2023, buhmann_caloclouds_2023, hashemi_deep_2024] or sensors [hashemi_ultra-high-resolution_2023], and possibly permutation invariant loss functions [tsan_particle_2021, di_bello_conditional_2022, kansal_evaluating_2023]. For normalizing flows, equivariance under permutation group action is more non-trivial. Proved by Köhler et al. [kohler_equivariant_2020], given a Flow-based model $F$ such that the set creation yields an exchangeable distribution, the update is permutation equivariant and invertible, and $p_{\theta}$ denotes the model likelihood, then $(F,-\log p_{\theta})$ is permutation equivariant.
@@
While maintaining the variable-length assumption, dropping the permutation equivariance would correspond to a sequential inductive bias of detector responses either layer-by-layer [hashemi_deep_2024, diefenbacher_l2lflows_2023], or hit-by-hit [lu_sarm_2021, liu_geometry-aware_2022, liu_generalizing_2023]. While autoregressive models show a better predictive capability [butter_jet_2023], the sequential nature of these models can be a disadvantage when it comes to high-dimensional computation, as the sampling is rather slow.
@@ 4.3 Out-Of-Distribution (OOD) Generation
DGMs for OOD and zero-shot learning is an exciting area of research that holds significant potential across various fields [anishchenko_novo_2021, freschlin_machine_2022, yeh_novo_2023, gainza_novo_2023, rajak_autonomous_2021, ravuri_skilful_2021, madani_large_2023, li_zero-knowledge_2023] including, but not limited to, drug Discovery, material design, and weather forecasting. Traditional methods for the synthetic generation of objects with enhanced or specific properties are often iterative and costly, requiring extensive manual work or heavy computational resources.
@@
In contrast, DGMs with zero-shot capability can deal with new scenarios that are not explicitly present in the training data, making them highly desirable in a wide range of problems. OOD generation of detector signatures is an emerging field in High Energy Physics (HEP) as well [paganini_calogan_2018, di_sipio_dijetgan_2019, howard_learning_2022, anderlini_generative_2023, liu_generalizing_2023, hashemi_deep_2024]. For example, for real PXD data at Belle II [abe_belle_2010], an important challenge in working is its reliance on actual experiments for gathering the necessary random triggers. Consequently, real luminosity and beam-parameter-dependent PXD background data that go beyond current experimental limits are unavailable, leaving us reliant on computationally demanding simulations. This underscores the critical need for a surrogate model that can effectively generalize to OOD luminosity regions. The current main challenge remains to be the optimization of DGMs, based on available information while avoiding overfitting, and the generalization to cases in which information is scarce or altogether absent, such as extrapolation to beam parameters, energies, luminosities, and geometries where there are no data.
@@
The first example of such OOD detector response generation is the CaloGAN paper [paganini_calogan_2018], where they very briefly show they can generate showers beyond the training incident energy conditions. DijetGAN [di_sipio_dijetgan_2019], also discusses extrapolation to new Beyond the Standard Model (BSM)-dependent OOD regions of dijet invariant mass. Anderlini et al. [anderlini_generative_2023] evaluates the uncertainty of GANs in new momentum regions through Background efficiency comparison. Their analysis shows that for Kaons, where the background efficiency does not decrease linearly in the OOD, high-momentum phase-space regions, their ensemble model fails to capture this behavior. Whereas, for Muons, due to the monotonic behavior of the background efficiency, the extrapolation shows better results.
@@
A very important inductive bias for OOD generation is the variable-length assumption, especially in the context of detector signature simulation, due to the inherently variable nature of the data, characterized by a fluctuating number of detector responses in individual events. DGMs with a variable-length inductive bias have a promising capability to generate sets/sequences of differing lengths, contingent on the inherent complexity or dynamism of the event being generated. This equips the model with better adaptability when faced with novel and unseen scenarios. Liu et al. [liu_geometry-aware_2022, liu_generalizing_2023], by using an ARM approach (length-independent), tries to extrapolate to unseen detector geometries where the extrapolation domain is calorimeter cell sizes. However, they show that the high granularity is a bottleneck to their approach. Hashemi et al. [hashemi_deep_2024] address the challenge of OOD simulation for high granular and sparse events (variable intra-event cardinality) by introducing YonedaVAE that can achieve both length and context extrapolation. Their model is trained on random trigger PXD background data from an early experiment of Belle II with a peak recorded luminosity of $1.42\times 10^{34}\text{cm}^{-2}\text{s}^{-1}$ and a mean occupancy of 0.06% ($\mathcal{O}(100)$ hit multiplicity). It was then tested on the previously unseen data from a later experiment of Belle II, which had nearly double the peak luminosity — $2.68\times 10^{34}\text{cm}^{-2}\text{s}^{-1}$ — and a mean occupancy of 0.32% ($\mathcal{O}(10^{5})$ hit multiplicity). The length extrapolation here was when the model had to infer how an event with a higher (than training) cardinality looks like. Context extrapolation means that the model has to produce the correct distribution of sparse cardinalities over all sensors, given only the maximum cardinality per event.
4.4 Anomaly Detection
Anomaly detection in high-energy physics is a large and growing field in itself (see for example [Kasieczka:2021xcg, Aarrestad:2021oeb] for reviews), but many of the proposed algorithms rely (at least in part) on DGMs. In general, we identify three main ideas that are currently explored.
The first one use the generative aspect of DGMs. The majority of these models are based on weakly supervised methods, especially the “Classification WithOut LAbels” (CWoLA) algorithm [Metodiev:2017vrx, Collins:2018epr, Collins:2019jip]. These methods require a signal-depleted background dataset, which can either be obtained in a data-driven way or from simulations by conditional generative models. Most of these anomaly detectors are based on Normalizing Flows [Hallin:2021wme, Raine:2022hht, Hallin:2022eoq, Mastandrea:2022vas, Golling:2022nkl, Sengupta:2023xqy, Golling:2023yjq, Finke:2023ltw, Bickendorf:2023nej, Bai:2023yyy] or diffusion models [Buhmann:2023acn, Sengupta:2023vtm]. There are also other approaches that use generative networks without the subsequent classifier step [Dillon:2019cqt, Dillon:2020quc, Fanelli:2022xwl].
The second group of anomaly detectors that use (parts of) DGMs is based on (V)AEs. This group uses the fact that the bottleneck architecture forces the encoder to learn a low-dimensional representation of “usual” data and anomalies can then be identified by a larger reconstruction error on the decoded output. It does therefore not rely on the generative nature of (V)AEs. This group constitutes the largest among all anomaly detectors that utilize generative architectures. In addition to only using autoencoders [Hajer:2018kqm, Heimel:2018mkt, Farina:2018fyg, Cerri:2018anq, Roy:2019jae, Blance:2019ibf, Amram:2020ykb, Cheng:2020dal, Pol:2020weg, vanBeekveld:2020txa, Bortolato:2021zic, Dillon:2021nxw, Finke:2021sdf, Atkinson:2021nlt, Govorkova:2021utb, Ostdiek:2021bem, Fraser:2021lxm, Herrero-Garcia:2021goa, Mikuni:2021nwn, Canelli:2021aps, Bradshaw:2022qev, Dillon:2022mkq, Golling:2023juz, Chekanov:2023uot, CMSECAL:2023fvz, Zhang:2023khv, Liu:2023djx], some further use components of GANs [Knapp:2020dde, Vaslin:2023lig] or Normalizing Flows [Park:2020pak, Jawahar:2021vyu, Buss:2022lxw].
The third group of anomaly detectors use DGMs that are at the same time density estimators. These can directly identify anomalous events by their small likelihood. Usually, these models only focus on the density estimation and anomaly detection aspect of the model [Nachman:2020lpy, Stein:2020rou, Caron:2021wmq, Verheyen:2022tov, Mikuni:2023tok, Das:2023bcj], but as was pointed out in [Krause:2023uww], they in principle have a double use: a single training of the DGM could be used for generating new samples and identifying anomalous events at no additional training costs, a promising new direction to optimize hardware resources.
@@ 5 Outlook: Challenges and Opportunities
Physics-Informed Generative Models. One of the most pressing challenges in the field of deep generative models for simulating detector signatures in particle physics is the incorporation of physics-based constraints or rules into the model architecture. While current methodologies often focus on data-driven approaches (permutation equivariance is a special case), they sometimes overlook the underlying physics and symmetries that govern the data. This could lead to models that are statistically accurate but physically implausible. Future work could focus on models that not only learn from data but also respect the governing symmetries [desai_symmetrygan_2022], thereby ensuring both statistical and physical fidelity in the generated signatures.
@@
Precision and Uncertainty Quantification. One of the less explored but critically important areas in the application of deep generative models for simulating high-granular detector signatures is the quantification of model uncertainty and the evaluation of precision [matchev_uncertainties_2022]. While these models can generate data that statistically resemble the target data, the question of how accurately they capture the underlying physical processes remains. Traditional metrics [krause_caloflow_2023-1, kansal_evaluating_2023, hashemi_ultra-high-resolution_2023] like log-likelihood or NN-based metrics provide some measure of the model’s performance but often fall short in capturing the nuances of physical validity and uncertainty. Moreover, the stochastic nature of these models introduces an inherent level of uncertainty in the predictions, which needs to be rigorously quantified to make the models truly useful. Future work could focus on the development of new evaluation metrics and techniques that not only assess the model’s ability to replicate known phenomena but also its reliability in predicting new high-granular detector signatures. This could involve Bayesian approaches [butter_generative_2023, das_how_2023, heimel_precision-machine_2023], ensemble methods [anderlini_generative_2023], or even hybrid models [dherin_morse_2023] to provide error bars along with predictions.
@@
Real-Time Optimization. The increasing complexity of particle physics experiments, characterized by high event rates and pile-up conditions, poses significant challenges for real-time data analysis and decision-making for ultra-high granular detectors. Surrogate models, as discussed in this paper, provide very affordable solutions; however, optimizing their real-time performance is still a bottleneck. One promising avenue is the use of Federated Learning (FL) [ratnayake_review_2023], which allows for decentralized optimization across multiple nodes, thereby leveraging the computational resources of the entire experimental setup. FL enables models to train on large real-time data sets and reduce biases associated with locally trained models. Meta-learning techniques [naik_meta-neural_1992] are also strong candidates to enable models to quickly adapt to new experimental data or detector conditions, reducing the time required for retraining. Recently, the MetaHEP [salamani_metahep_2023] project embarked on this path. Additionally, another promising avenue for future research lies in the direction of differentiable programming (DP) [adelmann_new_2022, dorigo_toward_2022, aehle_progress_2023, kagan_branches_2023]. Through DP, software becomes differentiable through automatic differentiation (AD) [baydin_automatic_2018], enabling efficient gradient computation to understand the influence of input variations on output predictions. Such gradient-based insights could be invaluable for various downstream tasks. Utilizing DP frameworks would allow particle physics simulation tools to be integrated into machine learning pipelines in an end-to-end manner, thereby facilitating joint optimization for enhanced computational efficiency.
@@
Quantum Generative Models. The advent of quantum computing offers a new avenue for tackling the computational challenges in particle physics simulations [di_meglio_quantum_2023]. Quantum Generative Models (QGM) could potentially revolutionize the field by providing exponential speed-ups for certain types of problems. However, the practical implementation of QGMs for particle physics is still in its infancy [bravo-prieto_style-based_2022, rehm_full_2023, rousselot_generative_2023, hoque_caloqvae_2023], and significant challenges related to error correction, qubit stability, and algorithmic design remain. Future research could focus on the development and validation of QGM architectures specifically tailored for detector signature simulations.
@@
Extrapolation Beyond Training Data. One of the limitations of current generative models is their ability to generalize beyond the scope of the training data. Given the high-dimensional nature of HEP data, the inevitability of sparse datasets arises due to limited sample sizes, leading us into zones with inadequate training data. If the model demonstrates strong extrapolation capabilities, it does not just fill gaps in areas completely lacking data; it also provides a more reliable interpretation of regions where data is minimal. This is particularly crucial in particle physics, where there is a need to explore uncharted territories of parameter space [howard_learning_2022], kinematic regions [butter_machine_2023, anderlini_generative_2023], luminosities [hashemi_deep_2024] or detector geometries [liu_generalizing_2023]. Developing models capable of reliable extrapolation is an open challenge and involves techniques like uncertainty quantification and the incorporation of prior physical knowledge and inductive bias into the model.
@@
Scalability and Ultra-High Granularity Challenges. The next generation of particle detectors will feature ultra-high granularities, as we defined in Section 3.2, leading to an explosion in the dimensionality of the data [apostolakis_detector_2022], for instance, the high Granularity Calorimeter (HGCAL) [magnan_hgcal_2017] at CMS [noauthor_phase-2_2017] with roughly $6$ million channels, the ultra-high resolution PXD at Belle II [abe_belle_2010] with more than $7.5$ million-pixel channels, or the EPICAL-2 [noauthor_results_nodate] electromagnetic calorimeter with $12.5$ million-pixel channels. Except for some efforts [hashemi_ultra-high_nodate, hashemi_deep_2024], current generative models are far from handling such ultra-high-dimensional data efficiently. Even these aforementioned efforts have not delved deeply into the issue of uncertainty quantification. This gap underscores the urgent need for future work that not only addresses the development of scalable algorithms and architectures that can handle ultra-high granularity detectors with irregular geometries but also rigorously quantifies the uncertainties to ensure the physical validity and reliability of the generated detector signatures.